Let X1, · · · , X20 be independent Poisson random variables with mean (print please)

1. (i) Use the pmf of Poisson distribution to compute P(X1 + · · · + X20 > 15).

(ii) Use the Markov inequality to obtain a bound on P(X1 + · · · + X20 > 15).

(iii) Use the central limit theorem to approximate P(X1 + · · · + X20 > 15).
Answer #1

Assuming lambda = 1

i)

Y = X1+X2+...+X20

Y follow poisson (20* lambda) = Pois(20)

P(Y > 15)

= 1- P(Y <= 15)

= 1- poisson.dist(15,20,1)

=

0.843487

ii) and iii)